OpenStudy (anonymous):

A compound made of two elements, iridium (Ir) and oxygen (O), was produced in a lab by heating iridium while exposed to air. The following data was collected: Mass of crucible: 38.26 g Mass of crucible and iridium: 39.52 g Mass of crucible and iridium oxide: 39.73 g Show, or explain, your calculations as you determine the empirical formula of the compound.

OpenStudy (anonymous):

ok do you have any idea as to how to start his?

OpenStudy (anonymous):

*this

OpenStudy (anonymous):

no.. not really

OpenStudy (anonymous):

well, start by subtracting the mass of the crucible from all your givens and tell me what you get.

OpenStudy (anonymous):

1.26 and 1.47

OpenStudy (anonymous):

ok, then tell me the molar masses of both iridium and oxygen

OpenStudy (anonymous):

16 and 192.22?

OpenStudy (anonymous):

ok now you need to calculate amount of each, namely Ir and O n(Ir)= 1,26/192,22 = 6,555 * 10^-3 mol n(O) = 0,21 / 16 = 0,013125 mol and now you need to see which has less amount and that is Ir and divide larger one with lesser one, in other words you divide amount of O with amount of Ir to see their ratio 0,013125/6,555*10^-3 = 2 so your compound is IrO2